If log(base b)X = 0.8 determine the value of log(baseb)b∜x
*can I just isolate the equation to get log(base b)x and sub in the value of 0.8?

Answer by lwsshak3(11628)

If log(base b)X = 0.8 determine the value of log(baseb)b∜x
**
logb(x)=0.8
logb(x^1/4)=(1/4)logb(x)=0.8/4=0.2